Factory Overhead Rates and Account Balance

Prostheses Industries operates two factories. The manufacturing operations of Factory 1 are machine intensive, while the manufacturing operations of Factory 2 are labor intensive. The company applies factory overhead to jobs on the basis of machine hours in Factory 1 and on the basis of direct labor hours in Factory 2. Estimated factory overhead costs, direct labor hours, and machine hours are as follows:

Factory 1 Factory 2
Estimated factory overhead cost for fiscal
year beginning August 1 $359,760 $655,200
Estimated direct labor hours for year 8,400
Estimated machine hours for year 14,990
Actual factory overhead costs for August $28,780 $56,910
Actual direct labor hours for August 760
Actual machine hours for August 1,170

a. Determine the factory overhead rate for Factory 1. Round your answer to the nearest cent.
$ per machine hour

b. Determine the factory overhead rate for Factory 2. Round your answer to the nearest cent.
$ per direct labor hour

c. Determine the factory overhead applied to production in each factory for August.

Factory 1 $
Factory 2 $

d. Determine the balances of the factory accounts for each factory as of August 31, and indicate whether the amounts represent overapplied or underapplied factory overhead.

Factory 1 $
Factory 2 $

e. Explain why Factory 1 uses machine hours to allocate factory overhead while Factory 2 uses direct labor hours.
Factory overhead should be allocated using a base that is related to (causes) the overhead costs incurred. Factory 1 has a   manufacturing operation, and Factory 2 has a   manufacturing operation. Thus, Factory 1 uses machine hours, and Factory 2 uses direct labor hours to allocate factory overhead.
Answer #1
a
Factory overhead rate for Factory 1 24 per machine hour =359760/14990
b
Factory overhead rate for Factory 2 78 per direct labor hour =655200/8400
c
Factory 1 28080 =1170*24
Factory 2 59280 =760*78
d
Factory 1 700 Underapplied =28780-28080
Factory 2 2370 Overapplied =59280-56910
e
Factory overhead should be allocated using a base that is related to (causes) the overhead costs incurred. Factory 1 has a machine-intensive manufacturing operation, and Factory 2 has a labor-intensive manufacturing operation. Thus, Factory 1 uses machine hours, and Factory 2 uses direct labor hours to allocate factory overhead.